### v1

# Ramsey goodness of stars and fans for the Hajós graph

Abstract: Given two graphs $G_{1}$ and $G_{2}$, the Ramsey number $R(G_{1},G_{2})$ denotes the smallest integer $N$ such that any red-blue coloring of the edges of $K_{N}$ contains either a red $G_{1}$ or a blue $G_{2}$. Let $G_{1}$ be a graph with chromatic number $\chi$ and chromatic surplus $s$, and let $G_{2}$ be a connected graph with $n$ vertices. The graph $G_{2}$ is said to be Ramsey-good for the graph $G_{1}$ (or simply $G_{1}$-good) if, for $n\geq s$,
$R(G_{1},G_{2})=(\chi-1)(n-1)+s.$
The $G_{1}$-good property has been extensively studied for star-like graphs when $G_{1}$ is a graph with $\chi(G_{1})\geq 3$, as seen in works by Burr-Faudree-Rousseau-Schelp (J. Graph Theory, 1983), Li-Rousseau (J. Graph Theory, 1996), Lin-Li-Dong (European J. Combin., 2010), Fox-He-Wigderson (Adv. Combin., 2023), and Liu-Li (J. Graph Theory, 2025), among others. However, all prior results require $G_{1}$ to have chromatic surplus $1$. In this paper, we extend this investigation to graphs with chromatic surplus 2 by considering the Hajós graph $H_{a}$. For a star $K_{1,n}$, we prove that $K_{1,n}$ is $H_{a}$-good if and only if $n$ is even. For a fan $F_{n}$ with $n\geq 111$, we prove that $F_{n}$ is $H_{a}$-good.
Keywords: Ramsey goodness, the Hajós graph, fan
AMS Subject Classification: 05C55, 05D10
1 Introduction
The Ramsey number is a bivariate function that assigns a positive integer $R(G_{1},G_{2})$ to every pair of simple graphs $G_{1}$ and $G_{2}$. It is defined as the smallest integer $N$ such that any graph $\Gamma$ on $N$ vertices contains $G_{1}$ as a subgraph, or its complement $\overline{\Gamma}$ contains $G_{2}$ as a subgraph.
The study of Ramsey numbers for sparse graphs has flourished since the 1970s. At that time, Burr [2] established a general lower bound for the Ramsey number of any pair of graphs. Suppose $G_{2}$ is a connected graph, and let $\chi(G_{1})$ and $s(G_{1})$ denote the chromatic number and the chromatic surplus of $G_{1}$, respectively, where the chromatic surplus refers to the smallest size of a color class over all proper $\chi(G_{1})$-colorings of $G_{1}$. Then, if $|G_{2}|\geq s(G_{1})$,
$R(G_{1},G_{2})\geq(\chi(G_{1})-1)(|G_{2}|-1)+s(G_{1}).$ (1)
When equality holds in (1), the graph $G_{2}$ is said to be $G_{1}$-good, or equivalently, $G_{2}$ is Ramsey-good for $G_{1}$. In the special case where $G_{1}$ is the complete graph $K_{k}$, such graphs $G_{2}$ are referred to as $k$-good by Burr and Erdős [3].
When $G_{1}$ is the complete $(k+1)$-partite graph $K_{1,m_{1},m_{2},\ldots,m_{k}}$ and $G_{2}$ is a star, Burr, Faudree, Rousseau, and Schelp [4] proved that for $m=\min\{m_{i}\mid i\in[k]\}$ and sufficiently large $n$,
$R(K_{1,m_{1},m_{2},\ldots,m_{k}},K_{1,n})=\begin{cases}k\cdot(n+m-2)+1&\text{if both }m\text{ and }n\text{ are even},\\ k\cdot(n+m-1)+1&\text{otherwise}.\end{cases}$
It follows that $K_{1,n}$ is $K_{1,m_{1},m_{2},\ldots,m_{k}}$-good only when $m_{1}=1$, or when $m_{1}=2$ and $n$ is even.
A frequently studied class of star-like graphs is the fan. A fan $F_{n}$ is the graph formed by $n$ triangles sharing a common vertex; see Figure 1. The concept of the fan was first introduced into extremal graph theory by Erdős, Füredi, Gould, and Gunderson [9], who studied its extremal graph and Turán number. For results on the Ramsey numbers of fans, see [5, 8, 15, 21].
Figure 1: The fan $F_{3}$ (left) and the Hajós graph (right)
The concept of a fan can be extended to that of a generalized fan. Given a graph $H$, let $nH$ denote the disjoint union of $n$ copies of $H$, and let $H_{1}+H_{2}$ denote the join of two disjoint graphs $H_{1}$ and $H_{2}$, obtained by adding all edges between $V(H_{1})$ and $V(H_{2})$. The graph $K_{1}+nH$ is called a generalized fan. In particular, if $H$ is $K_{1}$, then $K_{1}+nH$ is the star $K_{1,n}$; if $H$ is $K_{2}$, then $K_{1}+nH$ is the fan $F_{n}$.
Li and Rousseau [14] investigated the Ramsey-goodness of generalized fans. For any graphs $H$ and $G$, they proved that $K_{1}+nH$ is Ramsey-good for $K_{2}+G$ when $n$ is sufficiently large.
Observe that in any proper coloring of the graph $K_{2}+G$, there are two color classes containing only one vertex each. What happens if we require only one color class to have a single vertex? Let $K_{1}+K_{k}(m)$ denote the complete $(k+1)$-partite graph in which one partite set has size $1$ and each of the remaining $k$ partite sets has size $m$. Lin, Li, and Dong [16] showed that for $k\geq 2$, if $m$ is odd or if $n|H|$ is odd, then
$R(K_{1}+K_{k}(m),K_{1}+nH)=k(n|H|+m-1)+1.$
It follows that whenever $m\geq 2$, the graph $K_{1}+nH$ is not Ramsey-good for $K_{1}+K_{k}(m)$ in this setting.
Subsequently, without relying on the Erdős-Stone-Simonovits stability theorem, Chung and Lin [7] determined the exact value of $R(K_{1,m_{1},m_{2},\ldots,m_{k}},K_{1}+nH)$ for large $n$ in all cases. Let $m=\min\{m_{i}\mid i\in[k]\}$. Then
$r(K_{1,m_{1},m_{2},\ldots,m_{k}},K_{1}+nH)=\begin{cases}k\cdot(n|H|+m-2)+1&\text{if both }m\text{ and }n|H|\text{ are even},\\ k\cdot(n|H|+m-1)+1&\text{otherwise}.\end{cases}$
Another frequently studied star-like graph is the book $B_{k,n}$, which is the graph $K_{k}+(n-k)K_{1}$. Equivalently, $B_{k,n}$ can be viewed as the graph obtained by blowing up the center of the star $K_{1,n-k}$ into a clique $K_{k}$, while preserving the adjacency between each vertex of $K_{k}$ and the remaining vertices. For results concerning the Ramsey goodness of $B_{k,n}$, we refer the reader to [12, 19, 18]. For results on the Ramsey non-goodness of $B_{k,n}$, see [10, 11, 17].
It is easy to observe that in all the aforementioned works, when $G_{2}$ is a star or a star-like graph, the corresponding graph $G_{1}$ is always required to have chromatic surplus $1$. We extend these results by considering $G_{1}$ to be the Hajós graph $H_{a}$, which has chromatic surplus $2$. The Hajós graph, named after the Hungarian mathematician György Hajós, is a graph consisting of six vertices and nine edges. It is constructed by starting with a triangle and, for each of its edges, adding a new vertex and joining it to both endpoints of that edge; see Figure 1.
In this paper, we establish two main results concerning Ramsey goodness for the Hajós graph.
Our first result establishes that the star $K_{1,n}$ is $H_{a}$-good if and only if $n$ is even.
Theorem 1.
$R(H_{a},K_{1,n})=\begin{cases}2n+2&\text{for even }n\geq 2,\\ 2n+3&\text{for odd }n\geq 3.\end{cases}$
Our second result determines that $F_{n}$ is $H_{a}$-good for $n\geq 111$.
Theorem 2.
$R(H_{a},F_{n})=4n+2$ for $n\geq 111$.
We present the proofs of the two theorems in Section 2 and Section 3, respectively. The proof of Theorem 2 is more involved, and the lower bound $n\geq 111$ is not the best possible; rather, it arises from the limitations of our method. Therefore, identifying a smaller positive integer $n$ such that $F_{n}$ is $H_{a}$-good is a worthwhile direction for further research.
At the end of this section, we introduce some concepts and notation that will be used throughout. We use $[n]$ to denote the set $\{1,2,\ldots,n\}$. For a graph $G$ and a vertex subset $X\subseteq V(G)$, the subgraph induced by $X$ is denoted by $G[X]$, whose vertex set is $X$ and whose edge set consists of all edges in $G$ with both endpoints in $X$. For $u\in V(G)$ and $X\subseteq V(G)$, let $N_{X}(u)$ denote the set of neighbors of $u$ in $X$, and let $|N_{X}(u)|$ denote the number of such neighbors, also written as $d_{X}(u)$. When $X=V(G)$, we write $N_{X}(u)$ and $d_{X}(u)$ simply as $N(u)$ and $d(u)$, respectively. The minimum and maximum degrees of $G$ are denoted by $\delta(G)$ and $\Delta(G)$, respectively. Notation and terminology not explicitly defined follow Bondy and Murty [1].
2 Proof of Theorem 1
When $n$ is even, the lower bound can be derived from inequality (1):
$R(H_{a},K_{1,n})\geq(\chi(H_{a})-1)(|K_{1,n}|-1)+s(H_{a})=2n+2.$
When $n$ is odd, let $\ell=(n+1)/2$. Consider the graph $(\ell K_{2})+(\ell K_{2})$, which consists of two disjoint copies of a matching with $\ell$ edges, where the edges between the two copies form a complete bipartite graph. It is straightforward to verify that this graph does not contain the Hajós graph as a subgraph, and that the maximum degree in its complement is $n-1$. Therefore, $R(H_{a},K_{1,n})\geq 2n+3$ for odd $n$.
For the upper bound, we first consider the case $n=2$, namely, proving that $R(H_{a},K_{1,2})\leq 6$. For any graph $H$ of order $6$, if its complement $\overline{H}$ does not contain $K_{1,2}$ as a subgraph, then the edge set of $\overline{H}$ forms a matching. Consequently, $H$ must contain $K_{2,2,2}$ as a subgraph, which in turn contains $H_{a}$ as a subgraph.
Now suppose $n\geq 3$. Let $G$ be an arbitrary graph with $2n+2+\mathbf{1}_{\text{odd}}(n)$ vertices, where $\mathbf{1}_{\text{odd}}(n)$ is an indicator function: $\mathbf{1}_{\text{odd}}(n)=0$ if $n$ is even, and $\mathbf{1}_{\text{odd}}(n)=1$ if $n$ is odd. Suppose $\overline{G}$ contains no copy of $K_{1,n}$. We now prove that $G$ must contain the Hajós graph as a subgraph.
Since $\Delta(\overline{G})\leq n-1$, it follows that $\delta(G)\geq n+2+\mathbf{1}_{\text{odd}}(n)$. We proceed by considering three cases.
Case 1.
The graph $G$ does not contain $K_{4}$ as a subgraph.
By the classical result of Chvátal [6], we have $R(K_{3},K_{1,n})=2n+1$. Therefore, the graph $G$ must contain a triangle; without loss of generality, let its vertices be $u_{1},u_{2},u_{3}$. For $1\leq i<j\leq 3$, the minimum degree condition implies
$d(u_{i})+d(u_{j})\geq 2\delta(G)\geq 2n+4+2\times\mathbf{1}_{\text{odd}}(n)\geq|G|+2.$
Thus, vertices $u_{1}$ and $u_{2}$ share a common neighbor distinct from $u_{3}$, denoted $u_{4}$; similarly, $u_{1}$ and $u_{3}$ share a common neighbor $u_{5}\neq u_{2}$, and $u_{2}$ and $u_{3}$ share a common neighbor $u_{6}\neq u_{1}$. Since $G$ does not contain $K_{4}$ as a subgraph, the vertices $u_{4},u_{5},u_{6}$ must be pairwise distinct. Consequently, these six vertices induce a Hajós graph.
Case 2.
The graph $G$ contains $K_{5}-e$ as a subgraph.
Let us denote the five vertices of one such $K_{5}-e$ by $v_{1},v_{2},v_{3},v_{4},v_{5}$, and let the set of remaining vertices be $W$. Then we have $|W|=2n-3+\mathbf{1}_{\text{odd}}(n)$. Note that $K_{5}-e$ contains a clique of four vertices; without loss of generality, assume that $v_{1},v_{2},v_{3},v_{4}$ induce a $K_{4}$. By the minimum degree condition, for each $i\in[4]$, the vertex $v_{i}$ has at least $n-2+\mathbf{1}_{\text{odd}}(n)$ neighbors in $W$. Since $n\geq 3$, it follows that
$4(n-2+\mathbf{1}_{\text{odd}}(n))>2n-3+\mathbf{1}_{\text{odd}}(n)=|W|.$
Therefore, among $v_{1},v_{2},v_{3},v_{4}$, there must exist two vertices, say $v_{1}$ and $v_{2}$, which share a common neighbor in $W$, denoted by $v_{6}$. Because the vertex $v_{5}$ has at least two neighbors in $\{v_{2},v_{3},v_{4}\}$, we may assume without loss of generality that $v_{5}$ is adjacent to both $v_{2}$ and $v_{3}$. Thus, we obtain a Hajós graph, where the vertex subsets $\{v_{1},v_{2},v_{3}\}$, $\{v_{1},v_{2},v_{6}\}$, $\{v_{2},v_{3},v_{5}\}$, and $\{v_{1},v_{3},v_{4}\}$ each induce a triangle.
Case 3.
The graph $G$ does not contain $K_{5}-e$ as a subgraph, but contains $K_{4}$ as a subgraph.
Let $w_{1},w_{2},w_{3},w_{4}$ be the vertices of a $K_{4}$ in $G$.
We first consider the case when $n$ is even. Since $\delta(G)\geq n+2$, each vertex $w_{i}$ for $i\in[4]$ has at least $n-1$ neighbors in $V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$. As $|V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}|=2n-2$, there must exist two vertices, say $w_{1}$ and $w_{2}$, that share a common neighbor in $V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$, denoted by $w_{5}$. Since $G$ does not contain $K_{5}-e$ as a subgraph, neither $w_{3}$ nor $w_{4}$ is adjacent to $w_{5}$.
Let $X=V(G)\setminus\{w_{1},w_{2},w_{3},w_{4},w_{5}\}$. Then $|X|=2n-3$. By the minimum degree condition,
$d_{X}(w_{1})\geq n-2,d_{X}(w_{2})\geq n-2,d_{X}(w_{3})\geq n-1,\text{and }d_{X}(w_{4})\geq n-1.$
If $w_{1}$ and $w_{3}$ have a common neighbor in $X$, denoted by $w_{6}$, then we can construct a Hajós graph in which the vertex subsets $\{w_{1},w_{2},w_{3}\}$, $\{w_{1},w_{2},w_{5}\}$, $\{w_{2},w_{3},w_{4}\}$, and $\{w_{1},w_{3},w_{6}\}$ each induce a triangle. Hence, $N_{X}(w_{1})\cap N_{X}(w_{3})=\emptyset$. By symmetry, we have $N_{X}(w_{2})\cap N_{X}(w_{3})=\emptyset$.
Since $|X|=2n-3$, $d_{X}(w_{1})\geq n-2$ and $d_{X}(w_{3})\geq n-1$, the sets $N_{X}(w_{1})$ and $N_{X}(w_{3})$ form a partition of $X$, and we have $|N_{X}(w_{1})|=n-2$ and $|N_{X}(w_{3})|=n-1$. Given that $d_{X}(w_{2})\geq n-2$ and $N_{X}(w_{2})\cap N_{X}(w_{3})=\emptyset$, it follows that $N_{X}(w_{1})=N_{X}(w_{2})$. Therefore,
$N(w_{1})\cap N(w_{2})=N_{X}(w_{1})\cup\{w_{3},w_{4},w_{5}\}\text{ and }|N(w_{1})\cap N(w_{2})|=n+1.$
Since $n$ is even, by the classical result of Harary [13],
$R(K_{1,2},K_{1,n})=n+1=|N(w_{1})\cap N(w_{2})|.$
By assumption, $\overline{G}$ does not contain $K_{1,n}$ as a subgraph. Hence, the graph $K_{1,2}$ appears within the common neighborhood of $w_{1}$ and $w_{2}$, and together with $w_{1}$ and $w_{2}$ induces a subgraph isomorphic to $K_{5}-e$. This contradicts the assumption that $G$ does not contain $K_{5}-e$ as a subgraph.
Now consider the case where $n$ is odd. Let $Y=V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$. Since $\delta(G)\geq n+3$, each vertex $w_{i}$ for $i\in[4]$ has at least $n$ neighbors in $Y$. Given that $|Y|=2n-1$, any two vertices among $w_{1},w_{2},w_{3},w_{4}$ share at least one common neighbor in $Y$. Without loss of generality, assume that $w_{1}$ and $w_{2}$ share a common neighbor $w_{5}$ in $Y$, and $w_{2}$ and $w_{3}$ share a common neighbor $w_{6}$ in $Y$. As $G$ does not contain $K_{5}-e$ as a subgraph, vertices $w_{5}$ and $w_{6}$ must be distinct. Consequently, we obtain a Hajós graph in which the vertex subsets $\{w_{1},w_{2},w_{3}\}$, $\{w_{1},w_{2},w_{5}\}$, $\{w_{2},w_{3},w_{6}\}$, and $\{w_{1},w_{3},w_{4}\}$ each induce a triangle. ∎
3 Proof Theorem 2
The proof of Theorem 2 relies on the following key lemma, which was established by Zhao and the latter two authors of this paper [20].
Lemma 1.
$R(W_{4},F_{n})=4n+1$ for $n\geq 111$.
Now we begin the proof of Theorem 2.
Proof.
The lower bound follows directly from inequality (1):
$R(H_{a},F_{n})\geq(\chi(H_{a})-1)(|F_{n}|-1)+s(H_{a})=(3-1)(2n+1-1)+2=4n+2.$
To prove the upper bound, we proceed by contradiction. Assume that a graph $G$ with $24n+$ 2 vertices contains neither $H_{a}$ as a subgraph nor $F_{n}$ as a subgraph in $\overline{G}$. We will derive contradictions in the following two cases.
Case 1.
$\Delta(\overline{G})\geq 2n+2$
Let $u$ be a vertex in $\overline{G}$ with the maximum degree. Denote by $H^{{}^{\prime}}$ the subgraph of $\overline{G}$ induced by the set of vertices adjacent to $u$ in $\overline{G}$, i.e., $H^{{}^{\prime}}=\overline{G}[N_{\overline{G}}(u)]$. Let $M$ be a maximum matching in $H^{{}^{\prime}}$. Then $M$ must contain exactly $n-1$ edges. This is because if $M$ has at least $n$ edges, these $n$ edges together with $u$ would form an $F_{n}$ in $\overline{G}$, a contradiction. Conversely, if $M$ has at most $n-2$ edges, the number of vertices in $H^{{}^{\prime}}$ not covered by $M$ (i.e., the remaining vertices after removing the vertices of $M$) is at least $62n+2-2(n-2)=$ 6. These vertices form a complete graph in $G$, which clearly contains $H_{a}$ as a subgraph, leading to a contradiction.
Since $\Delta(\overline{G})\geq 2n+2$, it follows that $|V(H^{{}^{\prime}})\setminus V(M)|\geq 4$, and the vertices in $V(H^{{}^{\prime}})\setminus V(M)$ induce a complete subgraph in $G$.
- If $|V(H^{{}^{\prime}})\setminus V(M)|\geq 6$, it is evident that $G$ contains $H_{a}$ as a subgraph.
- If $|V(H^{{}^{\prime}})\setminus V(M)|=5$, consider any edge $y_{1}y_{2}\in M$. At least one of $y_{1}$ or $y_{2}$ is adjacent to at least four vertices in $V(H^{{}^{\prime}})\setminus V(M)$ in $G$. Otherwise, $H^{{}^{\prime}}$ would contain a matching larger than $M$, contradicting the maximality of $M$. Without loss of generality, assume $y_{1}$ is adjacent to at least four vertices in $V(H^{{}^{\prime}})\setminus V(M)$. Then the subgraph of $G$ induced by $(V(H^{{}^{\prime}})\setminus V(M))\cup\{y_{1}\}$ must contain $H_{a}$ as a subgraph, leading to a contradiction.
- If $|V(H^{{}^{\prime}})\setminus V(M)|=4$, consider any two edges $y_{1}y_{2}$ and $y_{3}y_{4}\in M$. For each edge, at least one endpoint must be adjacent to at least three vertices in $V(H^{{}^{\prime}})\setminus V(M)$ in $G$. Otherwise, $H^{{}^{\prime}}$ would contain a matching larger than $M$, again contradicting the maximality of $M$. Without loss of generality, assume $y_{1}$ and $y_{3}$ are each adjacent to at least three vertices in $V(H^{{}^{\prime}})\setminus V(M)$. Then the subgraph of $G$ induced by $(V(H^{{}^{\prime}})\setminus V(M))\cup\{y_{1},y_{3}\}$ must contain $H_{a}$ as a subgraph, leading to a contradiction.
Case 2.
$\delta(G)\geq 2n$
According to Lemma 1, the graph $G$ contains $W_{4}$ as a subgraph. Let the center of this $W_{4}$ be $u_{0}$, with the corresponding cycle $u_{1}u_{2}u_{3}u_{4}u_{1}$. We have the following claim.
Claim 1.
The edges $u_{1}u_{3},u_{2}u_{4}\not\in E(G)$.
Proof.
We only need to prove that $u_{1}u_{3}\not\in E(G)$; the other case can be proven similarly. Suppose $u_{1}u_{3}\in E(G)$. Since $\delta(G)\geq 2n$, the set $\{u_{1},u_{2},u_{3}\}$ must contain two vertices, say $u_{1}$ and $u_{3}$, that have a common neighbor in $V(G)\setminus V(W_{4})$. Otherwise, this would contradict the fact that $G$ has $24n+$ 2 vertices. Let $u_{5}$ be a common neighbor of $u_{1}$ and $u_{3}$ in $V(G)\setminus V(W_{4})$. Then the subgraph induced by the vertex set $\{u_{0},u_{1},u_{2},u_{3},u_{4},u_{5}\}$ in $G$ contains $H_{a}$ as a subgraph, a contradiction. ∎
If $u_{1}$ and $u_{2}$ have a common neighbor distinct from $u_{0}$, denoted by $u_{5}$, then the subgraph induced by the vertex set $\{u_{0},u_{1},u_{2},u_{3},u_{4},u_{5}\}$ contains $H_{a}$ as a subgraph. Therefore, the only common neighbor of $u_{1}$ and $u_{2}$ can be $u_{0}$. Furthermore, since $\delta(G)\geq 2n$, $u_{2}$ must have at least $12n-$ 1 neighbors in $V(G)\setminus N(u_{1})$. Similarly, $u_{4}$ must also have at least $12n-$ 1 neighbors in $V(G)\setminus N(u_{1})$. Hence, the number of common neighbors of $u_{2}$ and $u_{4}$ in $V(G)\setminus N(u_{1})$ is at least
$42(2n-1)-|V(G)\setminus N(u_{1})|\geq 2(2n-1)-(4n+2-2n)=2n-$ 4 .
We denote the set of common neighbors of vertices $u_{2}$ and $u_{4}$, excluding $u_{0}$, as $U_{1}$. Then, $|U_{1}|\geq 2n-4$. Similarly, let $U_{2}$ represent the set of common neighbors of vertices $u_{1}$ and $u_{3}$, excluding $u_{0}$. By symmetry, we have $|U_{2}|\geq 2n-4$. Since $u_{1}$ and $u_{2}$ have only one common neighbor, $u_{0}$, it follows that $U_{1}\cap U_{2}=\emptyset$. It is easy to verify that $u_{1},u_{3}\in U_{1}$ and $u_{2},u_{4}\in U_{2}$. To avoid the appearance of the graph $H_{a}$, for $i=1,3$ and $j=2,4$, the only common neighbor of $u_{i}$ and $u_{j}$ must be $u_{0}$. Therefore, the vertices $u_{1}$ and $u_{3}$ are isolated in the graph $G[U_{1}]$, and the vertices $u_{2}$ and $u_{4}$ are isolated in the graph $G[U_{2}]$. Furthermore, we have the following claim.
Claim 2.
Every connected component of the graph $G[U_{1}\setminus\{u_{1},u_{3}\}]$ is either an isolated vertex or a star.
Proof.
First, we prove that the graph $G[U_{1}\setminus\{u_{1},u_{3}\}]$ does not contain $P_{4}$ as a subgraph. We proceed by contradiction. Suppose $G[U_{1}\setminus\{u_{1},u_{3}\}]$ contains a $P_{4}$ as a subgraph, with vertices $x_{1},x_{2},x_{3},x_{4}$. The subgraph induced by the vertex set $\{u_{2},u_{4},x_{1},x_{2},x_{3},x_{4}\}$ in $G$ contains $H_{a}$ as a subgraph, which leads to a contradiction. Thus, each connected component of the graph $G[U_{1}\setminus\{u_{1},u_{3}\}]$ is either an isolated vertex, a star, or a triangle. Note that when a connected component is $K_{2}$, it is also considered a star.
Next, we only need to prove that the graph $G[U_{1}\setminus\{u_{1},u_{3}\}]$ does not contain a triangle as a subgraph. Again, we proceed by contradiction. Suppose $G[U_{1}\setminus\{u_{1},u_{3}\}]$ contains a triangle with vertices $x_{1},x_{2},x_{3}$. There must exist two vertices among $\{x_{1},x_{2},x_{3}\}$, say $x_{1}$ and $x_{3}$, that have a common neighbor in $V(G)\setminus\{x_{1},x_{2},x_{3},u_{2},u_{4}\}$, otherwise this would contradict the fact that the number of vertices in $G$ is $24n+$ 2. Let $x_{4}$ be a common neighbor of $x_{1}$ and $x_{3}$ in $V(G)\setminus\{x_{1},x_{2},x_{3},u_{2},u_{4}\}$. Then the subgraph induced by the vertex set $\{x_{1},x_{2},x_{3},x_{4},u_{2},u_{4}\}$ in $G$ contains $H_{a}$ as a subgraph, which leads to a contradiction. ∎
By symmetry, it can also be shown that each connected component of the graph $G[U_{2}\setminus\{u_{2},u_{4}\}]$ is either an isolated vertex or a star. Hence, the following claim holds.
Claim 3.
$|U_{1}|\leq 2n$ and $|U_{2}|\leq 2n$.
Proof.
We proceed by contradiction. Assume $|U_{1}|\geq 2n+1$ or $|U_{2}|\geq 2n+1$. By symmetry, we may assume $|U_{1}|\geq 2n+1$. It suffices to find a fan $F_{n}$ centered at $u_{1}$ in $\overline{G}$, leading to a contradiction.
By the previous claim, each connected component of the graph $G[U_{1}\setminus\{u_{1}\}]$ is either an isolated vertex or a star. Let the nontrivial (non-isolated) connected components of this graph be $C_{1},\ldots,C_{k}$. Note that in each connected component, at most one vertex has degree greater than $1$ in $G[U_{1}\setminus\{u_{1}\}]$.
When $k\geq 2$, for each $11\leq i\leq k-$ 1, the center vertex of $C_{i}$ and a leaf vertex of $C_{i+1}$ form an edge in $\overline{G}$. Similarly, the center vertex of $C_{k}$ and a leaf vertex of $C_{1}$ form an edge in $\overline{G}$. These $k$ edges form a matching in $\overline{G}$. The remaining vertices in $\overline{G}[U_{1}\setminus\{u_{1}\}]$ induce a complete subgraph, allowing us to find a matching of $n$ edges in $\overline{G}[U_{1}\setminus\{u_{1}\}]$. This gives a fan $F_{n}$ centered at $u_{1}$ in $\overline{G}$, a contradiction.
When $k=1$, the graph $G[U_{1}\setminus\{u_{1}\}]$ has only one nontrivial connected component, which must be a star. Let $v_{0}$ be the center of this star. Then $v_{0}u_{3}\in E(\overline{G})$. Since the remaining vertices in $\overline{G}[U_{1}\setminus\{u_{1}\}]$ induce a complete subgraph, $v_{0}u_{3}$ can be extended to a matching of $n$ edges in $\overline{G}[U_{1}\setminus\{u_{1}\}]$. This gives a fan $F_{n}$ centered at $u_{1}$ in $\overline{G}$, a contradiction.
When $k=0$, the graph $G[U_{1}\setminus\{u_{1}\}]$ is an empty graph, so $\overline{G}[U_{1}\setminus\{u_{1}\}]$ contains a matching of $n$ edges. This also gives a fan $F_{n}$ centered at $u_{1}$ in $\overline{G}$, a contradiction. ∎
We denote the set of vertices outside $U_{1}\cup U_{2}\cup\{u_{0}\}$ by $W$. Then,
$|W|\leq 4n+2-2(2n-4)-1=9.$
Furthermore, we partition $W$ into four subsets: $W_{1}$, $W_{2}$, $W_{3}$, and $W_{4}$. If a vertex in $W$ is adjacent to $u_{2}$ or $u_{4}$, it belongs to $W_{1}$. If a vertex in $W$ is adjacent to $u_{1}$ or $u_{3}$, it belongs to $W_{2}$. If a vertex $w\in W$ is not adjacent to any of $u_{1}$, $u_{2}$, $u_{3}$, or $u_{4}$, and the number of its neighbors in $U_{2}$ is at least the number of its neighbors in $U_{1}$, i.e., $d_{U_{2}}(w)\geq d_{U_{1}}(w)$, then it belongs to $W_{3}$. Otherwise, if $d_{U_{2}}(w)<d_{U_{1}}(w)$, it belongs to $W_{4}$. To avoid the appearance of the graph $H_{a}$, we have $W_{1}\cap W_{2}=\emptyset$. Consequently, $W_{1}$, $W_{2}$, $W_{3}$, and $W_{4}$ form a partition of $W$.
Note that the neighbors of $u_{2}$ must belong to $W_{1}\cup U_{1}\cup\{u_{0}\}$. Since $\delta(G)\geq 2n$, we have $|W_{1}|\geq 2n-1-|U_{1}|$. By symmetry, $|W_{2}|\geq 2n-1-|U_{2}|$. Hence, $|W_{3}|+|W_{4}|\leq(4n+2)-1-2(2n-1)=3$.
By the pigeonhole principle, either $|U_{1}|+|W_{1}|+|W_{3}|\geq 2n+1$, or $|U_{2}|+|W_{2}|+|W_{4}|\geq 2n+1$. Without loss of generality, assume the former holds. Next, we will show that if $G$ does not contain $H_{a}$ as a subgraph, then $\overline{G}[U_{1}\cup W_{1}\cup W_{3}]$ must contain a fan $F_{n}$ centered at $u_{1}$.
Note that every vertex in $(U_{1}\cup W_{1}\cup W_{3})\setminus\{u_{1},u_{3}\}$ is not adjacent to $u_{1}$ or $u_{3}$. Therefore, it suffices to find a matching of $n$ edges in $\overline{G}[(U_{1}\cup W_{1}\cup W_{3})\setminus\{u_{1}\}]$, which yields a fan $F_{n}$ centered at $u_{1}$ in $\overline{G}$.
From $W_{1}\cup W_{3}$, select $2n+1-|U_{1}|$ vertices, denoted by $w_{1},\ldots,w_{t}$, where $t=2n+1-|U_{1}|$. Since $|U_{1}|\geq 2n-4$, it follows that $t\leq 5$. We first find a matching $M_{1}$ in $\overline{G}[\{w_{1},\ldots,w_{t}\}\cup(U_{1}\setminus\{u_{1},u_{3}\})]$ such that each edge of $M_{1}$ has at least one endpoint in $\{w_{1},\ldots,w_{t}\}$ and $V(M_{1})$ includes as many vertices as possible from $\{w_{1},\ldots,w_{t}\}$. The following claim holds:
Claim 4.
At most one vertex in $\{w_{1},\ldots,w_{t}\}$ does not belong to $V(M_{1})$.
Proof.
Suppose at least three vertices in $\{w_{1},\ldots,w_{t}\}$ are not in $V(M_{1})$. Without loss of generality, assume $w_{1},w_{2},w_{3}\notin V(M_{1})$. It follows that $w_{1},w_{2},w_{3}$ form a triangle in $G$. To see this, if any edge is missing from $G$, say $w_{1}w_{2}\in E(\overline{G})$, we could add the edge $w_{1}w_{2}$ to $M_{1}$, forming a new matching that includes more vertices from $\{w_{1},\ldots,w_{t}\}$, contradicting the choice of $M_{1}$. Additionally, by the choice of $M_{1}$, $w_{1}$, $w_{2}$, and $w_{3}$ must each be adjacent to every vertex in $U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))$. Since $t\leq 5$, at most two vertices of $V(M_{1})$ are in $U_{1}$. Thus, $|U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))|\geq(2n-4)-(2+2)=2n-8$. Selecting any three vertices from $U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))$, together with $w_{1},w_{2},w_{3}$, induces a complete multipartite graph $K_{1,1,1,3}$ in $G$, which contains $H_{a}$ as a subgraph, a contradiction.
If exactly two vertices in $\{w_{1},\ldots,w_{t}\}$ are not in $V(M_{1})$, assume $w_{1},w_{2}\notin V(M_{1})$. Then $w_{1}w_{2}\in E(G)$ by the choice of $M_{1}$. Moreover, $w_{1}$ and $w_{2}$ must be adjacent to every vertex in $U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))$. Since at most three vertices in $V(M_{1})$ are in $U_{1}$, we have $|U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))|\geq|U_{1}|-5\geq 2n-9$.
First, consider $w_{1},w_{2}\in W_{1}$. If $w_{1}$ and $w_{2}$ are both adjacent to either $u_{2}$ or $u_{4}$, without loss of generality, assume $w_{1}u_{2},w_{2}u_{2}\in E(G)$. Selecting any three vertices from $U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))$, together with $w_{1}$, $w_{2}$, and $u_{2}$, induces a complete multipartite graph $K_{1,1,1,3}$ in $G$, which contains $H_{a}$ as a subgraph, leading to a contradiction. If $w_{1}$ and $w_{2}$ are adjacent to different vertices in $\{u_{2},u_{4}\}$, assume $w_{1}u_{2},w_{2}u_{4}\in E(G)$. Selecting any two vertices $u_{x},u_{y}$ from $U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))$, the four sets $\{w_{1},w_{2},u_{x}\}$, $\{w_{1},w_{2},u_{y}\}$, $\{w_{1},u_{2},u_{y}\}$, and $\{w_{2},u_{4},u_{y}\}$ each induce a triangle. Thus, $G$ contains $H_{a}$ as a subgraph, a contradiction.
Next, consider $w_{1},w_{2}\in W_{3}$. By the definition of $W_{3}$, $w_{1}$ and $w_{2}$ each have at least $92n-$ 9 neighbors in $U_{2}$. Thus, $w_{1}$ and $w_{2}$ share at least $182(2n-9)-|U_{2}|\geq 2n-$ 18 common neighbors in $U_{2}$, where this inequality follows from Claim 3.
If each common neighbor of $w_{1}$ and $w_{2}$ in $U_{2}$ has degree at least $2$ in $G[U_{2}]$, note that each connected component in $G[U_{2}]$ has at most one vertex of degree at least $2$. Hence, $G[U_{2}]$ must have at least $182n-$ 18 components, each containing a vertex of degree at least $2$, contradicting $|U_{2}|\leq 2n$. Thus, there exists a common neighbor $u_{z}$ of $w_{1}$ and $w_{2}$ in $U_{2}$ with degree at most $1$ in $G[U_{2}]$. Then the vertex $u_{z}$ has at least $112n-2-|W|\geq 2n-$ 11 neighbors in $U_{1}$. Consequently, $u_{z}$, $w_{1}$, and $w_{2}$ share at least $(2n-11)+(|U_{1}|-5)-|U_{1}|\geq 2n-16$ common neighbors in $U_{1}$. Selecting three such common neighbors together with $u_{z},w_{1},w_{2}$ induces a complete multipartite graph $K_{1,1,1,3}$, which contains $H_{a}$ as a subgraph, a contradiction.
If $w_{1}$ and $w_{2}$ belong to $W_{1}$ and $W_{3}$, respectively, let $w_{1}\in W_{1}$ and $w_{2}\in W_{3}$. Then $w_{1}$ must be adjacent to either $u_{2}$ or $u_{4}$. By symmetry, assume $w_{1}u_{2}\in E(G)$. According to the selection of $W_{3}$, $w_{2}$ has at least $92n-$ 9 neighbors in $U_{2}$. Therefore, $w_{2}$ must have a neighbor $u_{z}$ in $U_{2}$ whose degree in $G[U_{2}]$ is at most 1; otherwise, this would contradict $|U_{2}|\leq 2n$. Consequently, the number of neighbors of $u_{z}$ in $U_{1}$ is at least $112n-2-|W|\geq 2n-$ 11. Thus, the number of common neighbors of $u_{z}$, $w_{1}$, and $w_{2}$ in $U_{1}$ is at least $(2n-11)+(|U_{1}|-5)-|U_{1}|\geq 2n-16$. Selecting two common neighbors, together with $u_{z}$, $w_{1}$, $w_{2}$, and $u_{2}$, forms a subgraph in $G$ that contains $H_{a}$ as a subgraph, leading to a contradiction. ∎
Based on Claim 2, every connected component in the graph $G[U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))]$ is either an isolated vertex or a star. Denote the connected components of this graph as $D_{1},\ldots,D_{k}$. Note that each connected component contains at most one vertex with degree greater than $1$ in $G[U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))]$. Let $H=G[(U_{1}\cup\{w_{1},\ldots,w_{t}\})\setminus(\{u_{1}\}\cup V(M_{1}))]$. We establish the following claim.
Claim 5.
There exists a matching $M_{2}$ in $\overline{H}$ that covers all vertices in $H$ with degree greater than $1$.
Proof.
When $k\geq 2$, for $i\in[k]$ and $D_{k+1}:=D_{1}$, if $D_{i}$ contains a vertex with degree greater than $1$, then this vertex has no edge to a vertex with degree at most $1$ in $D_{i+1}$. That is, there exists an edge between these two vertices in $\overline{H}$. Add this edge to $M_{2}$. If exactly one vertex in $\{w_{1},\ldots,w_{t}\}$ is not in $V(M_{1})$, assume without loss of generality that $w_{1}\notin V(M_{1})$. Then $w_{1}u_{3}\in E(\overline{G})$, and we add the edge $w_{1}u_{3}$ to $M_{2}$. If $\{w_{1},\ldots,w_{t}\}\subseteq V(M_{1})$, the edge $w_{1}u_{3}$ does not need to be included in $M_{2}$. Clearly, $M_{2}$ covers all vertices in $H$ with degree greater than $1$.
When $k=1$, the graph $G[U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))]$ has only one connected component, which must be a star. Let the center of this star be $v_{0}$, and let two of its leaves be $v_{1}$ and $v_{2}$.
If $\{w_{1},\ldots,w_{t}\}\subseteq V(M_{1})$, add the edge $v_{0}u_{3}$ to $M_{2}$. Since $v_{0}$ is the only vertex in $H$ with degree greater than $1$, $M_{2}$ covers all such vertices.
If exactly one vertex in $\{w_{1},\ldots,w_{t}\}$ is not in $V(M_{1})$, assume without loss of generality that $w_{1}\notin V(M_{1})$. By the choice of $M_{1}$, $w_{1}$ is adjacent to all vertices in $U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))$, including $v_{0}$, $v_{1}$, and $v_{2}$. This implies that $w_{1}$ has at least $112n-$ 11 neighbors in $U_{1}$.
If $w_{1}\in W_{3}$, by the definition of $W_{3}$, $w_{1}$ has at least $112n-$ 11 neighbors in $U_{2}$. Since $v_{2}$ has exactly one neighbor in $U_{1}$, it must have at least $112n-2-|W|\geq 2n-$ 11 neighbors in $U_{2}$. Thus, $w_{1}$ and $v_{2}$ share at least 2($222(2n-11)-|U_{2}|\geq 2n-$ 22 common neighbors in $U_{2}$. Let $u^{\prime}$ be one such common neighbor. The sets $\{w_{1},v_{0},v_{1}\}$, $\{w_{1},v_{0},v_{2}\}$, $\{w_{1},v_{2},u^{\prime}\}$, and $\{u_{2},v_{0},v_{2}\}$ each induce a triangle. Therefore, $G$ contains $H_{a}$ as a subgraph, a contradiction.
If $w_{1}\in W_{1}$, then either $w_{1}u_{2}\in E(G)$ or $w_{1}u_{4}\in E(G)$. By symmetry, assume $w_{1}u_{2}\in E(G)$. The sets $\{w_{1},v_{0},v_{1}\}$, $\{w_{1},v_{0},v_{2}\}$, $\{w_{1},v_{2},u_{2}\}$, and $\{u_{4},v_{0},v_{2}\}$ each induce a triangle. Hence, $G$ contains $H_{a}$ as a subgraph, a contradiction. ∎
In the subgraph $G[(U_{1}\cup\{w_{1},\ldots,w_{t}\})\setminus(\{u_{1}\}\cup V(M_{1})\cup V(M_{2}))]$, each vertex has degree at most $1$, and this subgraph contains at least $2(n-|M_{1}|-|M_{2}|)$ vertices. Below, we demonstrate that $n-|M_{1}|-|M_{2}|\geq 2$.
Since in the graph $G[U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))]$, at most one-third of the vertices have degree greater than $1$, the vertex set $V(M_{2})$ can contain at most two-thirds of this portion. Therefore, the remaining vertices number at least $|U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))|/3\geq(2n-11)/3$. Combining this with $n\geq 111$, the subgraph $G[(U_{1}\cup\{w_{1},\ldots,w_{t}\})\setminus(\{u_{1}\}\cup V(M_{1})\cup V(M_{2}))]$ has at least $(222-11)/3\geq 70$ vertices.
Hence, in the graph $\overline{G}[(U_{1}\cup\{w_{1},\ldots,w_{t}\})\setminus(\{u_{1}\}\cup V(M_{1})\cup V(M_{2}))]$, we can identify a matching consisting of $n-|M_{1}|-|M_{2}|$ edges, denoted by $M_{3}$. In $\overline{G}$, the edge set $M_{1}\cup M_{2}\cup M_{3}$ forms a matching of $n$ edges, and every vertex in this matching is adjacent to the vertex $u_{1}$. Consequently, $\overline{G}$ contains $F_{n}$ as a subgraph, leading to a final contradiction. ∎
Acknowledgements
Y. Zhang was partially supported by the National Natural Science Foundation of China (NSFC) under Grant No. 11601527 and by the Natural Science Foundation of Hebei Province under Grant No. A2023205045.
Data Availability Statement
No data was used or generated in this research.
Conflict of interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@
 Ramsey goodness of stars and fans for the Hajós graph
-Abstract: Given two graphs $G_{1}$ and $G_{2}$, the Ramsey number $R(G_{1},G_{2})$ denotes the smallest integer $N$ such that any red-blue coloring of the edges of $K_{N}$ contains either a red $G_{1}$ or a blue $G_{2}$.
+Abstract
+Given two graphs $G_{1}$ and $G_{2}$, the Ramsey number $R(G_{1},G_{2})$ denotes the smallest integer $N$ such that any red-blue coloring of the edges of $K_{N}$ contains either a red $G_{1}$ or a blue $G_{2}$.
 Let $G_{1}$ be a graph with chromatic number $\chi$ and chromatic surplus $s$, and let $G_{2}$ be a connected graph with $n$ vertices.
@@
 For a fan $F_{n}$ with $n\geq 111$, we prove that $F_{n}$ is $H_{a}$-good.
-Keywords: Ramsey goodness, the Hajós graph, fan
-AMS Subject Classification: 05C55, 05D10
+keywords:
+Ramsey goodness, the Hajós graph, fan
 1 Introduction
@@
 The study of Ramsey numbers for sparse graphs has flourished since the 1970s.
-At that time, Burr [2] established a general lower bound for the Ramsey number of any pair of graphs.
+At that time, Burr (1981) established a general lower bound for the Ramsey number of any pair of graphs.
 Suppose $G_{2}$ is a connected graph, and let $\chi(G_{1})$ and $s(G_{1})$ denote the chromatic number and the chromatic surplus of $G_{1}$, respectively, where the chromatic surplus refers to the smallest size of a color class over all proper $\chi(G_{1})$-colorings of $G_{1}$.
@@
 When equality holds in (1), the graph $G_{2}$ is said to be $G_{1}$-good, or equivalently, $G_{2}$ is Ramsey-good for $G_{1}$.
-In the special case where $G_{1}$ is the complete graph $K_{k}$, such graphs $G_{2}$ are referred to as $k$-good by Burr and Erdős [3].
-When $G_{1}$ is the complete $(k+1)$-partite graph $K_{1,m_{1},m_{2},\ldots,m_{k}}$ and $G_{2}$ is a star, Burr, Faudree, Rousseau, and Schelp [4] proved that for $m=\min\{m_{i}\mid i\in[k]\}$ and sufficiently large $n$,
+In the special case where $G_{1}$ is the complete graph $K_{k}$, such graphs $G_{2}$ are referred to as $k$-good by Burr and Erdős (1983).
+When $G_{1}$ is the complete $(k+1)$-partite graph $K_{1,m_{1},m_{2},\ldots,m_{k}}$ and $G_{2}$ is a star, Burr et al. (1983) proved that for $m=\min\{m_{i}\mid i\in[k]\}$ and sufficiently large $n$,
 $R(K_{1,m_{1},m_{2},\ldots,m_{k}},K_{1,n})=\begin{cases}k\cdot(n+m-2)+1&\text{if both }m\text{ and }n\text{ are even},\\ k\cdot(n+m-1)+1&\text{otherwise}.\end{cases}$
@@
 A fan $F_{n}$ is the graph formed by $n$ triangles sharing a common vertex; see Figure 1.
-The concept of the fan was first introduced into extremal graph theory by Erdős, Füredi, Gould, and Gunderson [9], who studied its extremal graph and Turán number.
-For results on the Ramsey numbers of fans, see [5, 8, 15, 21].
+The concept of the fan was first introduced into extremal graph theory by Erdős et al.
+(1995), who studied its extremal graph and Turán number.
+For results on the Ramsey numbers of fans, see Chen et al.
+(2021); Dvořák and Metrebian (2023); Lin and Li (2009); Zhang et al.
+(2015).
 Figure 1: The fan $F_{3}$ (left) and the Hajós graph (right)
@@
 In particular, if $H$ is $K_{1}$, then $K_{1}+nH$ is the star $K_{1,n}$; if $H$ is $K_{2}$, then $K_{1}+nH$ is the fan $F_{n}$.
-Li and Rousseau [14] investigated the Ramsey-goodness of generalized fans. For any graphs $H$ and $G$, they proved that $K_{1}+nH$ is Ramsey-good for $K_{2}+G$ when $n$ is sufficiently large.
+Li and Rousseau (1996) investigated the Ramsey-goodness of generalized fans. For any graphs $H$ and $G$, they proved that $K_{1}+nH$ is Ramsey-good for $K_{2}+G$ when $n$ is sufficiently large.
 Observe that in any proper coloring of the graph $K_{2}+G$, there are two color classes containing only one vertex each.
@@
 Let $K_{1}+K_{k}(m)$ denote the complete $(k+1)$-partite graph in which one partite set has size $1$ and each of the remaining $k$ partite sets has size $m$.
-Lin, Li, and Dong [16] showed that for $k\geq 2$, if $m$ is odd or if $n|H|$ is odd, then
+Lin et al.
+(2010) showed that for $k\geq 2$, if $m$ is odd or if $n|H|$ is odd, then
 $R(K_{1}+K_{k}(m),K_{1}+nH)=k(n|H|+m-1)+1.$
 It follows that whenever $m\geq 2$, the graph $K_{1}+nH$ is not Ramsey-good for $K_{1}+K_{k}(m)$ in this setting.
-Subsequently, without relying on the Erdős-Stone-Simonovits stability theorem, Chung and Lin [7] determined the exact value of $R(K_{1,m_{1},m_{2},\ldots,m_{k}},K_{1}+nH)$ for large $n$ in all cases.
+Subsequently, without relying on the Erdős-Stone-Simonovits stability theorem, Chung and Lin (2025) determined the exact value of $R(K_{1,m_{1},m_{2},\ldots,m_{k}},K_{1}+nH)$ for large $n$ in all cases.
 Let $m=\min\{m_{i}\mid i\in[k]\}$.
@@
 Equivalently, $B_{k,n}$ can be viewed as the graph obtained by blowing up the center of the star $K_{1,n-k}$ into a clique $K_{k}$, while preserving the adjacency between each vertex of $K_{k}$ and the remaining vertices.
-For results concerning the Ramsey goodness of $B_{k,n}$, we refer the reader to [12, 19, 18].
-For results on the Ramsey non-goodness of $B_{k,n}$, see [10, 11, 17].
+For results concerning the Ramsey goodness of $B_{k,n}$, we refer the reader to Fox et al.
+(2023); Nikiforov and Rousseau (2004); Liu and Li (2025).
+For results on the Ramsey non-goodness of $B_{k,n}$, see Fan and Lin (2023); Fan et al.
+(2024); Lin and Liu (2021).
 It is easy to observe that in all the aforementioned works, when $G_{2}$ is a star or a star-like graph, the corresponding graph $G_{1}$ is always required to have chromatic surplus $1$.
@@
 The minimum and maximum degrees of $G$ are denoted by $\delta(G)$ and $\Delta(G)$, respectively.
-Notation and terminology not explicitly defined follow Bondy and Murty [1].
+Notation and terminology not explicitly defined follow Bondy and Murty (2008).
 2 Proof of Theorem 1
@@
 Consider the graph $(\ell K_{2})+(\ell K_{2})$, which consists of two disjoint copies of a matching with $\ell$ edges, where the edges between the two copies form a complete bipartite graph.
-It is straightforward to verify that this graph does not contain the Hajós graph as a subgraph, and that the maximum degree in its complement is $n-1$.
+It is straightforward to verify that this graph does not contain the Hajós graph as a subgraph, and that the maximum degree in its complement is n1n-1.
 Therefore, $R(H_{a},K_{1,n})\geq 2n+3$ for odd $n$.
@@
 The graph $G$ does not contain $K_{4}$ as a subgraph.
-By the classical result of Chvátal [6], we have $R(K_{3},K_{1,n})=2n+1$.
+By the classical result of Chvátal (1977), we have $R(K_{3},K_{1,n})=2n+1$.
 Therefore, the graph $G$ must contain a triangle; without loss of generality, let its vertices be $u_{1},u_{2},u_{3}$.
@@
 We first consider the case when $n$ is even.
-Since $\delta(G)\geq n+2$, each vertex $w_{i}$ for $i\in[4]$ has at least $n-1$ neighbors in $V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$.
+Since $\delta(G)\geq n+2$, each vertex $w_{i}$ for $i\in[4]$ has at least n1n-1 neighbors in $V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$.
 As $|V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}|=2n-2$, there must exist two vertices, say $w_{1}$ and $w_{2}$, that share a common neighbor in $V(G)\setminus\{w_{1},w_{2},w_{3},w_{4}\}$, denoted by $w_{5}$.
 Since $G$ does not contain $K_{5}-e$ as a subgraph, neither $w_{3}$ nor $w_{4}$ is adjacent to $w_{5}$.
-Let $X=V(G)\setminus\{w_{1},w_{2},w_{3},w_{4},w_{5}\}$. Then $|X|=2n-3$. By the minimum degree condition,
+Let $X=V(G)\setminus\{w_{1},w_{2},w_{3},w_{4},w_{5}\}$. Then |X|=2n3|X|=2n-3. By the minimum degree condition,
 $d_{X}(w_{1})\geq n-2,d_{X}(w_{2})\geq n-2,d_{X}(w_{3})\geq n-1,\text{and }d_{X}(w_{4})\geq n-1.$
@@
 By symmetry, we have $N_{X}(w_{2})\cap N_{X}(w_{3})=\emptyset$.
-Since $|X|=2n-3$, $d_{X}(w_{1})\geq n-2$ and $d_{X}(w_{3})\geq n-1$, the sets $N_{X}(w_{1})$ and $N_{X}(w_{3})$ form a partition of $X$, and we have $|N_{X}(w_{1})|=n-2$ and $|N_{X}(w_{3})|=n-1$.
+Since |X|=2n3|X|=2n-3, $d_{X}(w_{1})\geq n-2$ and $d_{X}(w_{3})\geq n-1$, the sets $N_{X}(w_{1})$ and $N_{X}(w_{3})$ form a partition of $X$, and we have $|N_{X}(w_{1})|=n-2$ and $|N_{X}(w_{3})|=n-1$.
 Given that $d_{X}(w_{2})\geq n-2$ and $N_{X}(w_{2})\cap N_{X}(w_{3})=\emptyset$, it follows that $N_{X}(w_{1})=N_{X}(w_{2})$.
@@
 $N(w_{1})\cap N(w_{2})=N_{X}(w_{1})\cup\{w_{3},w_{4},w_{5}\}\text{ and }|N(w_{1})\cap N(w_{2})|=n+1.$
-Since $n$ is even, by the classical result of Harary [13],
+Since $n$ is even, by the classical result of Harary (1972),
 $R(K_{1,2},K_{1,n})=n+1=|N(w_{1})\cap N(w_{2})|.$
@@
 Since $\delta(G)\geq n+3$, each vertex $w_{i}$ for $i\in[4]$ has at least $n$ neighbors in $Y$.
-Given that $|Y|=2n-1$, any two vertices among $w_{1},w_{2},w_{3},w_{4}$ share at least one common neighbor in $Y$.
+Given that |Y|=2n1|Y|=2n-1, any two vertices among $w_{1},w_{2},w_{3},w_{4}$ share at least one common neighbor in $Y$.
 Without loss of generality, assume that $w_{1}$ and $w_{2}$ share a common neighbor $w_{5}$ in $Y$, and $w_{2}$ and $w_{3}$ share a common neighbor $w_{6}$ in $Y$.
@@
 ∎
-3 Proof Theorem 2
-The proof of Theorem 2 relies on the following key lemma, which was established by Zhao and the latter two authors of this paper [20].
+3 Proof of Theorem 2
+The proof of Theorem 2 relies on the following key lemma, which was established by Zeng et al. (2025).
 Lemma 1.
@@
 Now we begin the proof of Theorem 2.
-Proof.
+of Theorem 2.
 The lower bound follows directly from inequality (1):
@@
 To prove the upper bound, we proceed by contradiction.
-Assume that a graph $G$ with $24n+$ 2 vertices contains neither $H_{a}$ as a subgraph nor $F_{n}$ as a subgraph in $\overline{G}$.
+Assume that a graph $G$ with $4n+2$ vertices contains neither $H_{a}$ as a subgraph nor $F_{n}$ as a subgraph in $\overline{G}$.
 We will derive contradictions in the following two cases.
@@
 Let $M$ be a maximum matching in $H^{{}^{\prime}}$.
-Then $M$ must contain exactly $n-1$ edges.
+Then $M$ must contain exactly n1n-1 edges.
 This is because if $M$ has at least $n$ edges, these $n$ edges together with $u$ would form an $F_{n}$ in $\overline{G}$, a contradiction.
-Conversely, if $M$ has at most $n-2$ edges, the number of vertices in $H^{{}^{\prime}}$ not covered by $M$ (i.e., the remaining vertices after removing the vertices of $M$) is at least $62n+2-2(n-2)=$ 6.
+Conversely, if $M$ has at most n2n-2 edges, the number of vertices in $H^{{}^{\prime}}$ not covered by $M$ (i.e., the remaining vertices after removing the vertices of $M$) is at least 2n+22(n2)=62n+2-2(n-2)=6.
 These vertices form a complete graph in $G$, which clearly contains $H_{a}$ as a subgraph, leading to a contradiction.
@@
 Since $\delta(G)\geq 2n$, the set $\{u_{1},u_{2},u_{3}\}$ must contain two vertices, say $u_{1}$ and $u_{3}$, that have a common neighbor in $V(G)\setminus V(W_{4})$.
-Otherwise, this would contradict the fact that $G$ has $24n+$ 2 vertices.
+Otherwise, this would contradict the fact that $G$ has $4n+2$ vertices.
 Let $u_{5}$ be a common neighbor of $u_{1}$ and $u_{3}$ in $V(G)\setminus V(W_{4})$.
@@
 Therefore, the only common neighbor of $u_{1}$ and $u_{2}$ can be $u_{0}$.
-Furthermore, since $\delta(G)\geq 2n$, $u_{2}$ must have at least $12n-$ 1 neighbors in $V(G)\setminus N(u_{1})$.
-Similarly, $u_{4}$ must also have at least $12n-$ 1 neighbors in $V(G)\setminus N(u_{1})$.
+Furthermore, since $\delta(G)\geq 2n$, $u_{2}$ must have at least 2n12n-1 neighbors in $V(G)\setminus N(u_{1})$.
+Similarly, $u_{4}$ must also have at least 2n12n-1 neighbors in $V(G)\setminus N(u_{1})$.
 Hence, the number of common neighbors of $u_{2}$ and $u_{4}$ in $V(G)\setminus N(u_{1})$ is at least
-$42(2n-1)-|V(G)\setminus N(u_{1})|\geq 2(2n-1)-(4n+2-2n)=2n-$ 4 .
+$2(2n-1)-|V(G)\setminus N(u_{1})|\geq 2(2n-1)-(4n+2-2n)=2n-4.$
 We denote the set of common neighbors of vertices $u_{2}$ and $u_{4}$, excluding $u_{0}$, as $U_{1}$.
@@
 Suppose $G[U_{1}\setminus\{u_{1},u_{3}\}]$ contains a triangle with vertices $x_{1},x_{2},x_{3}$.
-There must exist two vertices among $\{x_{1},x_{2},x_{3}\}$, say $x_{1}$ and $x_{3}$, that have a common neighbor in $V(G)\setminus\{x_{1},x_{2},x_{3},u_{2},u_{4}\}$, otherwise this would contradict the fact that the number of vertices in $G$ is $24n+$ 2.
+There must exist two vertices among $\{x_{1},x_{2},x_{3}\}$, say $x_{1}$ and $x_{3}$, that have a common neighbor in $V(G)\setminus\{x_{1},x_{2},x_{3},u_{2},u_{4}\}$, otherwise this would contradict the fact that the number of vertices in $G$ is $4n+2$.
 Let $x_{4}$ be a common neighbor of $x_{1}$ and $x_{3}$ in $V(G)\setminus\{x_{1},x_{2},x_{3},u_{2},u_{4}\}$.
@@
 Note that in each connected component, at most one vertex has degree greater than $1$ in $G[U_{1}\setminus\{u_{1}\}]$.
-When $k\geq 2$, for each $11\leq i\leq k-$ 1, the center vertex of $C_{i}$ and a leaf vertex of $C_{i+1}$ form an edge in $\overline{G}$.
+When $k\geq 2$, for each $1\leq i\leq k-1$, the center vertex of $C_{i}$ and a leaf vertex of $C_{i+1}$ form an edge in $\overline{G}$.
 Similarly, the center vertex of $C_{k}$ and a leaf vertex of $C_{1}$ form an edge in $\overline{G}$.
@@
 Next, consider $w_{1},w_{2}\in W_{3}$.
-By the definition of $W_{3}$, $w_{1}$ and $w_{2}$ each have at least $92n-$ 9 neighbors in $U_{2}$.
-Thus, $w_{1}$ and $w_{2}$ share at least $182(2n-9)-|U_{2}|\geq 2n-$ 18 common neighbors in $U_{2}$, where this inequality follows from Claim 3.
+By the definition of $W_{3}$, $w_{1}$ and $w_{2}$ each have at least 2n92n-9 neighbors in $U_{2}$.
+Thus, $w_{1}$ and $w_{2}$ share at least $2(2n-9)-|U_{2}|\geq 2n-18$ common neighbors in $U_{2}$, where this inequality follows from Claim 3.
 If each common neighbor of $w_{1}$ and $w_{2}$ in $U_{2}$ has degree at least $2$ in $G[U_{2}]$, note that each connected component in $G[U_{2}]$ has at most one vertex of degree at least $2$.
-Hence, $G[U_{2}]$ must have at least $182n-$ 18 components, each containing a vertex of degree at least $2$, contradicting $|U_{2}|\leq 2n$.
+Hence, $G[U_{2}]$ must have at least 2n182n-18 components, each containing a vertex of degree at least $2$, contradicting $|U_{2}|\leq 2n$.
 Thus, there exists a common neighbor $u_{z}$ of $w_{1}$ and $w_{2}$ in $U_{2}$ with degree at most $1$ in $G[U_{2}]$.
-Then the vertex $u_{z}$ has at least $112n-2-|W|\geq 2n-$ 11 neighbors in $U_{1}$.
+Then the vertex $u_{z}$ has at least $2n-2-|W|\geq 2n-11$ neighbors in $U_{1}$.
 Consequently, $u_{z}$, $w_{1}$, and $w_{2}$ share at least $(2n-11)+(|U_{1}|-5)-|U_{1}|\geq 2n-16$ common neighbors in $U_{1}$.
@@
 By symmetry, assume $w_{1}u_{2}\in E(G)$.
-According to the selection of $W_{3}$, $w_{2}$ has at least $92n-$ 9 neighbors in $U_{2}$.
+According to the selection of $W_{3}$, $w_{2}$ has at least 2n92n-9 neighbors in $U_{2}$.
 Therefore, $w_{2}$ must have a neighbor $u_{z}$ in $U_{2}$ whose degree in $G[U_{2}]$ is at most 1; otherwise, this would contradict $|U_{2}|\leq 2n$.
-Consequently, the number of neighbors of $u_{z}$ in $U_{1}$ is at least $112n-2-|W|\geq 2n-$ 11.
+Consequently, the number of neighbors of $u_{z}$ in $U_{1}$ is at least $2n-2-|W|\geq 2n-11$.
 Thus, the number of common neighbors of $u_{z}$, $w_{1}$, and $w_{2}$ in $U_{1}$ is at least $(2n-11)+(|U_{1}|-5)-|U_{1}|\geq 2n-16$.
@@
 By the choice of $M_{1}$, $w_{1}$ is adjacent to all vertices in $U_{1}\setminus(\{u_{1},u_{3}\}\cup V(M_{1}))$, including $v_{0}$, $v_{1}$, and $v_{2}$.
-This implies that $w_{1}$ has at least $112n-$ 11 neighbors in $U_{1}$.
-If $w_{1}\in W_{3}$, by the definition of $W_{3}$, $w_{1}$ has at least $112n-$ 11 neighbors in $U_{2}$.
-Since $v_{2}$ has exactly one neighbor in $U_{1}$, it must have at least $112n-2-|W|\geq 2n-$ 11 neighbors in $U_{2}$.
-Thus, $w_{1}$ and $v_{2}$ share at least 2($222(2n-11)-|U_{2}|\geq 2n-$ 22 common neighbors in $U_{2}$.
+This implies that $w_{1}$ has at least 2n112n-11 neighbors in $U_{1}$.
+If $w_{1}\in W_{3}$, by the definition of $W_{3}$, $w_{1}$ has at least 2n112n-11 neighbors in $U_{2}$.
+Since $v_{2}$ has exactly one neighbor in $U_{1}$, it must have at least $2n-2-|W|\geq 2n-11$ neighbors in $U_{2}$.
+Thus, $w_{1}$ and $v_{2}$ share at least $2(2n-11)-|U_{2}|\geq 2n-22$ common neighbors in $U_{2}$.
 Let $u^{\prime}$ be one such common neighbor.
@@
 ∎
-Acknowledgements
+Acknowledgements.
+The authors are grateful to the anonymous referee for helpful remarks.
 Y.